If a wire made up of unknown material is 25.0 m long with a diameter of 0.100 mm and has a resistance of 77.7Ω at 20.0°C, What is its resistance at 186 °C?


Round your answer to 2 decimal places


1
Expert's answer
2021-05-16T17:52:14-0400
"R=\rho\frac{l}{0.25\pi d^2}\\77.7=\frac{25}{0.25\pi (1\cdot10^{-4})^2}\rho\\\rho=2.44\cdot10^{-8}\Omega m"

It is Gold!


"R'=77.7(1+(0.0034)(186-20))=121.6\Omega"